The following estimated regression equation was developed relating yearly income (y in $1000s) of 30 individuals with their age (x1) and their gender (x2) (0 if male and 1 if female). Ŷ = 30 + .7x1 + 3x2 Also provided are SST = 1200 and SSE = 384. The test statistic for testing the significance of the model is
Answer #2

$$ \begin{aligned} &\text { SSR }=\text { SST - SSE } \\ &=1200-384 \\ &=816 \\ &R^{2}=\frac{S S R}{S S T}=\frac{816}{1200}=0.68 \end{aligned} $$

answered by: fisher
Answer #1

The test statistic for testing the significance of the model is =MSR/MSE =(SSR/(k))/(SSE/(n-k-1))

=((1200-384)/2)/(384/(30-2-1))=28.6875